Four times the sum of three and a number is 3. What is the number?
I have no clue how to set this up!

Answer by wilft1(217)

the sum of 2 numbers means your adding them, four times that will be...
4(3+x)=3
lets solve this
4 (3+x) =3
12 + 4x = 3
subtract 12 from both sides
4x = -9
divide both sides by 4
x = -2.25
plug that in and check your answer
4 ( 3 + (-2.25)) = 3
4 (0.75) = 3
3 = 3